Hello, I am a student studying abroad, currently i have to work on some assignments my professor gave me to solve. Since my professor doesnt have the best grasp on the english language i am having trouble understanding what he requires me to do.

I know its the holidays, but i could really use some help here ! Happy New Year to you all

Someone decided to deposit 100 € in a bank that pays an effective interest rate of 6% every six months. After 2,5 years and at the end of the following periods of six months he decided to make additional deposits of 40 € each, except at the end of the final period of the 6 months when he drew 90 € to buy bonds. Taking these data into account find out the accumulated capital at the end of 12 years time.
Answer #1

FV =PV* (1+r)^n

R= 6%/2=3%

N = 12*2=24

Accumulated capital= (€100*(1.03^24))+((40*(1.03^23))-(40*(1.03^5))-90 = 203.28+(78.94- 46.37)-90 = 203.28+32.57-90 = €145.85